# PromptFix: You Prompt and We Fix the Photo

### v1

@@ 5.2 Quantitative and Qualitative Results
Figure 3 illustrates the visual comparison across all selected baseline models. In colorization, our PromptFix produces the most visually accurate and vibrant results, closely resembling the ground truth. For watermark removal task, it effectively restores the original image without introducing artifacts, outperforming MGIE [19] and other methods. In desnowing and low-light enhancement, PromptFix achieves clearer and more natural outputs, significantly reducing noise and enhancing visibility. Additionally, in super-resolution, PromptFix demonstrates remarkable clarity and accurate, preserving fine details and surpassing all comparative methods. In dehazing, although PromptFix’s performance is visually comparable to image restoration experts PromptIR [51] and AirNet [38], PromptFix outperforms the recent stable-diffusion-based method Diff-plugin [46], achieving a clean, sharp appearance, and closely matching the ground truth.
5.3 Ablation Study
Effectiveness of High-frequency Guidance Sampling. The High-frequency Guidance Sampling (HGS) method is introduced to balance fidelity and quality. To validate the effectiveness of HGS, we conduct qualitative experiments. As depicted in Figure 5, in a low-light scenario, the model aims to enhance the visibility (quality) of the input image while preserving its original textual details (fidelity). For baselines leveraging stable-diffusion as a generative prior, the strong compression capability of VAE also brings the issue of spatial information loss, as demonstrated by InstructDiff [21], MGIE [19], and Diff-Plugin [46] in Figure 5. This issue is independent of the model’s ability to effectively follow instructions. As shown by the variant “Ours w/o HGS,” our method significantly enhances low-light images compared to the three baselines, yet still fails to retain small-scale textual structures. By incorporating HGS, as seen in “Ours,” the proposed framework delivers a high-fidelity solution that also meets the low-light enhancement instruction.
@@ 5.3 Ablation Study
VLM-guided blind restoration. We utilize LLaVA [43] to generate auxiliary prompts and leave the instruction prompt empty. This approach enables users to input an image without needing to provide instructions for its restoration. We evaluate the model’s performance on such blind restoration tasks, including low-light enhancement, desnowing, and dehazing. As shown in Table 3, our model achieves performance comparable to four baselines, showcasing minimal perceptual differences from the ground truth and superior zero-shot capabilities.
@@
Multi-task processing. Although PromptFix is not explicitly trained to handle multiple low-level tasks simultaneously within the same image, it demonstrates the capability for multi-task processing. We construct the validation dataset with 200 images, and each image contains 3 restoration tasks such as colorization, watermark removal, low-light enhancement, desnow, dehazy, and super-resolution. We benchmarked PromptFix against AirNet and PromptIR—both generalist image restoration methods—as well as InstructP2P and InstructDiff, which are instruction-driven diffusion methods. As shown in Table 3, PromptFix outperforms these baselines, achieving superior image quality, structural similarity, and minimal perceptual differences from the ground truth, as evidenced by competitive PSNR, SSIM, and LPIPS scores, along with a higher ManIQA score indicating visually pleasing and high-quality results. Conversely, while methods like InstructP2P and InstructDiff perform well in specific metrics, they do not match the overall balanced performance of PromptFix. These results indicate the robustness and versatility of PromptFix.
@@ 6 Conclusion
We present PromptFix, a novel diffusion-based model along with a large-scale visual-instruction training dataset, designed to benefit instruction-guided low-level image processing. PromptFix effectively addresses challenges related to spatial information loss and degradation adaptation by high-frequency guidance sampling and a VLM-based auxiliary prompt module. These mechanisms improve the model’s performance in the instruction-based image-processing paradigm. Extensive experiment results demonstrate PromptFix’s advanced capabilities of generating accurate and high-quality images. In addition to the improvement in conventional metrics, we observe that PromptFix is also effective at processing multi-task processing and achieving blind restoration in low-light enhancement, desnowing, and dehazing.
@@ Appendix A Appendix
A.1 Datasets
Object removal and creation: We collect a dataset with 300,000 image pairs based on the GQA dataset [30]. For each object in all images, we segment the object using the Segment Anything Model [35] with the bounding box as the prompt, and then utilize LaMa [62] to inpaint the object. For each object, we generate instructions based on its relation with other objects in the same image. We support object removal with raw image input, images with bounding box prompts, and images with points prompts; additionally, we support object creation with images with bounding box prompts and images with points prompts. For testing purposes, we select 308 pairs of images from the complete dataset.
Image dehazing: We use a combination of synthetic datasets (RESIDE [37], SRRS [11]) and real-world datasets (Dense-Haze [1], O-Haze [2]) for training, comprising 100 pairs of real-world images and 102,230 pairs of indoor and outdoor synthetic images from ITS [37], OTS [37], SOTS outdoor [37], and nyuhaze500 [37]. For testing purposes, we randomly select 300 pairs of images from the complete dataset.
Colorization: We utilize a subset of Laion-5b [57] comprising 46,465 images and generate grayscale images. For testing purposes, we randomly select 300 pairs of images from the dataset.
Super resolution: We use a combination of subsets from Laion-5b [57], Flickr30K [50], HQ-50K [67], Flickr1M, and Pexels100K, which together contain 99,603 images. All images in Laion-5b [57], Flickr30K [50], and HQ-50K [67] are downsampled using scales of 4, 6, 8, 12, and 16, resulting in a total of 288,015 image pairs, with approximately 200,000 pairs downsampled at scales of 12 and 16. For the Flickr1M and Pexels100K datasets, containing 197,973 images, they are downsampled to six specific resolutions of the long side of images from 128 to 384, forming a dataset with 315,000 image pairs. For testing purposes, we randomly select 300 pairs of images from the complete dataset.
Low light enhancement: We use a combination of LOL [66], SID [9], SMID [8], and SDSD [64], which collectively contain 47,139 pairs of images. For testing purposes, we select 336 pairs of images from the complete dataset.
Snow removal: We use a combination of SRRS [11], CSD [12], RVSD [10], and Snow100K [47], which collectively contain 60,879 pairs of images. For testing purposes, we randomly select 300 pairs of images from the complete dataset.
Watermark removal: We use a combination of CLWD [44] and LOGO30K [16], which collectively contain 124,805 pairs of images. For testing purposes, we randomly select 300 pairs of images from the complete dataset.
@@ A.2 Limitations
Without relying on user-input instructions, our PromptFix achieves blind restoration for low-level enhancement, dehazing, and desnowing. However, we observe that this approach occasionally results in out-of-conditioned image control, where the model performs text-to-image generation based on the auxiliary prompt rather than image processing. Although blind restoration is a feature of our VLM-based Auxiliary Prompt Module, for known degradations, we recommend providing user-custom instructions to specify the restoration.
@@
Figure 7: Exponential Decay Weight Functions
@@ A.3 More Implementation Details
Weight of Fidelity Constraint: The time-scale weight, which is determined by the time step of the diffusion model, increases the noise in the latent variables as the time step becomes larger. Consequently, the fidelity constraint should be decayed accordingly.
@@ A.4 Efficiency analysis
We conduct an efficiency analysis for PromptFix and other recent baseline models, including Diff-plugin and InstructDiff, by profiling the floating-point operations (FLOPs) of their respective diffusion models. As demonstrated in the table below, our method exhibits comparable efficiency to these advanced models. This analysis reveals the capability of PromptFix to maintain high computational efficiency while delivering superior performance in various image-processing tasks.
@@
Diff-Plugin 1.1
InstructDiff 0.8
@@ A.5 More Results
General PromptFix Results. We present additional image processing results in Figures 9-16, encompassing tasks including watermark removal, colorization, low-light enhancement, super-resolution, dehazing, desnowing, inpainting, and object removal via bounding boxes. These results are generated by our PromptFix model, utilizing a all-in-one pretrained weight and guided by user-customized instructions.
@@
Figure 8: Visual results for Mutli-task processing
Figure 9: More results for watermark removal
Figure 10: More results for image colorization
Figure 11: More results for low-light enhancement
Figure 12: More results for super resolution
Figure 13: More results for image dehazing
Figure 14: More results for desnow
Figure 15: More results for image inpainting
Figure 16: More results for object removal by bounding box
Figure 17: Visual results for VLM-guided blind restoration

### v2

@@ 5.3 Ablation Study
VLM-guided blind restoration. We utilize LLaVA [41] to generate auxiliary prompts and leave the instruction prompt empty. This approach enables users to input an image without the need to provide instructions for its restoration. We evaluate the model’s performance on such blind restoration tasks, including low-light enhancement, desnowing, and dehazing. As shown in Table 3, our model achieves performance comparable to four baselines, showing minimal perceptual differences from the ground truth and superior zero-shot capabilities.
@@
Multi-task processing. Although PromptFix is not explicitly trained to handle multiple low-level tasks simultaneously within the same image, it demonstrates the capability for multi-task processing. We construct the validation dataset with 200 images, and each image contains 3 restoration tasks such as colorization, watermark removal, low-light enhancement, desnowing, dehazing, and super-resolution. We benchmarked PromptFix against AirNet and PromptIR, both generalist image restoration methods, as well as InstructP2P and InstructDiff, which are instruction-driven diffusion methods. As shown in Table 3, PromptFix outperforms these baselines, achieving superior image quality, structural similarity, and minimal perceptual differences from the ground truth, as evidenced by competitive PSNR, SSIM, and LPIPS scores, along with a higher ManIQA score indicating visually pleasing and high-quality results. Conversely, while methods like InstructP2P and InstructDiff perform well in specific metrics, they do not match the overall balanced performance of PromptFix. These results indicate the robustness and versatility of PromptFix.
Different types of instruction prompts. We verify PromptFix’s generalization to various human instructions in Table 5, by conducting ablation comparisons with three types of prompts: instructions used during training and out-of-training human instructions with fewer than 20 words and with 40-70 words. PromptFix’s performance slightly declines with out-of-training instructions, but the change is negligible. It indicates that PromptFix is robust for instructions under 20 words, which is generally sufficient for low-level processing tasks. We observe a performance drop with longer instructions, possibly due to the long-tail effect of instruction length in the training data. Although low-level processing tasks usually don’t require long instructions, addressing this issue by augmenting the dataset with longer instructions could be a direction for future work.
Table 4: Quantitative study on HGS and Auxiliary Prompting. $\mathcal{F}(\cdot)$ and $\mathcal{S}(\cdot)$ represent the type of high-frequency operators used in HGS to guide sampling.
[…]
- - 0.2068 0.6487
- 0.1707 0.6300
- 0.1795 0.6195
- 0.1990 0.5856
0.1600 0.6274
Table 5: Ablation Study on Different Types of Instruction Prompt. $\mathbb{A}$: instructions used during training; $\mathbb{B}$: out-of-training human instructions with fewer than 20 words; $\mathbb{C}$: out-of-training human instructions with 40-70 words.
[…]
$\mathbb{A}$ 0.1600 0.6274
PromptFix $\mathbb{B}$ 0.1639 0.6258
$\mathbb{C}$ 0.1823 0.5958
@@ 6 Conclusion
We present PromptFix, a novel diffusion-based model along with a large-scale visual-instruction training dataset, designed to benefit instruction-guided low-level image processing. PromptFix effectively addresses challenges related to spatial information loss and degradation adaptation by high-frequency guidance sampling and a VLM-based auxiliary prompt module. These mechanisms improve the model’s performance in the instruction-based image-processing paradigm. Extensive experiment results demonstrate PromptFix’s advanced capabilities of generating accurate and high quality images. In addition to the improvement in terms of conventional metrics, we observe that PromptFix is also effective at processing multi-task processing and achieving blind restoration in low-light enhancement, desnowing, and dehazing.
@@ Appendix A Appendix
A.1 PromptFix Dataset
The PromptFix dataset contains 1,013,320 triplet instances among 7 different tasks containing object removal, image dehazing, colorization, debluring, low light enhancement, snow removal and watermark removal. Each instance contains an input image, a processed image, an instruction, and an auxiliary prompt generated by InternVL2 [13] which is only available for tasks except object removal. The dataset is available at https://huggingface.co/datasets/yeates/PromptfixData.
Object removal: Mulan [65] provides multi-layer image decomposition annotations, inpainting occluded areas and isolating instances in a scene on separate RGBA layers. We combine the background layer and $n$ foreground object layers to obtain the ground truth image and the background layer and $n+1$ foreground object layers to obtain the input image, where $n\in\mathbb{N}$. Different object layers will be placed onto the background image based on their foreground-background relationships.
Image dehazing: We use a combination of synthetic datasets (RESIDE [35], SRRS [11]) and real-world datasets (Dense-Haze [1], O-Haze [2]) for training, comprising 100 pairs of real-world images and 102,230 pairs of indoor and outdoor synthetic images from ITS [35], OTS [35], SOTS outdoor [35], and nyuhaze500 [35].
Colorization: We utilize a subset of Laion-5b [58] and Flickr comprising 317,225 images and generate grayscale images.
Deblurring: We use a combination of GoPro [47], HIDE [59], LFDOF [56], RealBlur [54], TextOCR [60], Wider-Face [69] and a subset of CelebA [46], which collectively contain 70,600 pairs of images.
Low light enhancement: We use a combination of LOL [68], SID [9], SMID [8], SDSD [66], and construct a set of low light synthetic data based on Pexels and Flickr by reducing brightness and adding noise, which collectively contains 212,618 pairs of images.
Snow removal: We use a combination of SRRS [11], CSD [12], RVSD [10], and Snow100K [45], which collectively contains 99,177 pairs of images.
Watermark removal: We use a combination of CLWD [42] and LOGO30K [16], which collectively contain 124,805 pairs of images.
@@ A.2 Limitations
Without relying on user input instructions, our PromptFix achieves blind restoration for low-level enhancement, dehazing, and desnowing. However, we observe that this approach occasionally results in out-of-conditioned image control, where the model performs text-to-image generation based on the auxiliary prompt rather than image processing. Although blind restoration is a feature of our VLM-based Auxiliary Prompt Module, for known degradations, we recommend providing user-custom instructions to specify the restoration.
@@
Figure 7: Exponential Decay Weight Functions.
@@ A.3 More Implementation Details
Weight of Fidelity Constraint: The time-scale weight, which is determined by the time step of the diffusion model, decreases as the time step increases, due to the increase in noise. Consequently, the fidelity constraint should be decayed accordingly.
@@ A.4 Efficiency analysis
We conduct an efficiency analysis for PromptFix and other recent baseline models, including Diff-plugin [44] and InstructDiff [21], by profiling the floating-point operations (FLOPs) of their respective diffusion models. As demonstrated in the table below, our method exhibits comparable efficiency to these advanced models. This analysis reveals the capability of PromptFix to maintain high computational efficiency while delivering superior performance in various image-processing tasks.
@@
Diff-Plugin [44] 1.1
InstructDiff [21] 0.8
@@ A.5 More Results
General PromptFix Results. We present additional image processing results in Figures 9-15, encompassing tasks including watermark removal, colorization, low-light enhancement, super-resolution, dehazing, desnowing, inpainting, and object removal via bounding boxes. These results are generated by our PromptFix model, utilizing an all-in-one pre-trained weight and guided by user-customized instructions.
@@
Figure 8: Visual results for Mutli-task processing.
Figure 9: More results for watermark removal.
Figure 10: More results for image colorization.
Figure 11: More results for low-light enhancement.
Figure 12: More results for image dehazing.
Figure 13: More results for desnowing.
Figure 14: More results for image inpainting.
Figure 15: More results for object removal by a bounding box.
Figure 16: Visual results for VLM-guided blind restoration.